# Computation of the semiclassical outflux emerging from a collapsing spherical null shell

### v1

@@ Abstract
We consider a minimally coupled, massless quantum scalar field $\hat{\Phi}$ propagating in the background geometry of a four-dimensional black hole formed by the collapse of a spherical thin null shell, with a Minkowski interior and a Schwarzschild exterior. The field is taken in the natural “in” vacuum state, namely, the quantum state in which no excitations arrive from past null infinity. Within the semiclassical framework, we analyze the vacuum polarization $\left\langle\hat{\Phi}^{2}\right\rangle_{\text{ren}}$ and the energy outflux density $\left\langle\hat{T}_{uu}\right\rangle_{\text{ren}}$ (where $u$ is the standard null Eddington coordinate) just outside the shell. Using the point-splitting method, we derive closed-form analytical expressions for both these semiclassical quantities. In particular, our result for $\left\langle\hat{T}_{uu}\right\rangle_{\text{ren}}$ reveals that it vanishes like $(1-2M_{0}/r_{0})^{2}$ as the shell collapses toward the event horizon, where $M_{0}$ is the shell’s mass and $r_{0}$ is the value of the area coordinate $r$ at the evaluation point. This confirms that, along a late-time outgoing null geodesic (i.e., one that emerges from the shell very close to the event horizon and propagates toward future null infinity), the outflux gradually evolves (from virtually zero) up to its final Hawking-radiation value while the geodesic traverses the strong-field region (rather than the Hawking-radiation outflux being emitted entirely from the collapsing shell, which would lead to significant backreaction effects).
@@ Introduction.
Where does this Hawking-radiation energy come from? This has been a long-standing question (ChenUnruhWuYeom:2018, ). It is quite obvious that it emerges out of the strong-field region, but it is less clear exactly from what part of that region it originates. A widely accepted view (see, e.g., Ref. (ChenUnruhWuYeom:2018, )), to which we shall refer as option (i), is that the outgoing radiation starts from zero at the event horizon (EH) and gradually develops over a fairly broad strong-field region, until the area coordinate $r$ reaches a few times the BH mass $M$. However, there has also been an alternative viewpoint (option (ii)) suggesting that this energy outflux actually emerges directly from the collapsing body itself, and propagates in a conserved (or approximately conserved) manner through the surrounding strong-field vacuum region until it arrives FNI as Hawking radiation (see e.g. (Davies:1976, ; KawaiMatsuoYokokura:2013, ; Ho:2015, ; KawaiYokokura:2016, ; Ho:2016, ; MaskalaniecSikorski:2024, )).
For simplicity, let us consider a spherically-symmetric collapsing object, surrounded by a vacuum region whose geometry is described by the Schwarzschild metric  [^1] We use general-relativistic units $c=G=1$. $ds^{2}=-f\,dt^{2}+f^{-1}dr^{2}+r^{2}d\Omega^{2}$, where $f\equiv 1-2M/r$ and $d\Omega^{2}\equiv d\theta^{2}+\sin^{2}\theta\,d\varphi^{2}$. We shall focus here on the domain $r>2M$ of this Schwarzschild geometry. We define the two null Eddington coordinates $u,v$ by $v\equiv t+r_{*}$ and $u\equiv t-r_{*}$, where $r_{*}\equiv r+2M\ln\left(r/2M-1\right)$ (in Minkowski, $r_{*}=r$). The energy outflux density is then represented by $\left\langle\hat{T}_{uu}\right\rangle_{\text{ren}}$, that is, the outgoing null component of the renormalized stress-energy tensor (RSET) $\left\langle\hat{T}_{\mu\nu}\right\rangle_{\text{ren}}(x)$.
@@
Evidently, this quantity diverges on approaching the EH, which would potentially lead to drastic backreaction effects – such as entirely halting the collapse process (KawaiMatsuoYokokura:2013, ; Mersini:2014(HH), ; Mersini:2014(Unruh), ; Ho:2015, ; KawaiYokokura:2016, ; Ho:2016, ; MaskalaniecSikorski:2024, ) and/or preventing the EH formation (for various scenarios of “horizon avoidance” in semiclassical collapse see, e.g., (KawaiMatsuoYokokura:2013, ; Mersini:2014(HH), ; Mersini:2014(Unruh), ; Ho:2015, ; KawaiYokokura:2016, ; Ho:2016, ; BMT:2016, ; BMT:2017a, ; BMT:2017b, ; MaskalaniecSikorski:2024, )).
@@
We hereafter model the collapse process by a thin null massive shell, with an empty, flat interior (more details below). For simplicity we consider a minimally-coupled massless scalar field $\Phi(x)$. Several authors (see, e.g., Refs. (AndersonSiahmazginClarkFabbri:2020, ; SiahmazgiAndersonClarkFabbri:2021, ) and references within) have already elaborated on this model, but so far, to our knowledge, a concrete result for $\left\langle\hat{T}_{uu}\right\rangle_{\text{ren}}$ in the strong-field region has not been obtained. We shall derive a simple, explicit, analytical expression for $\left\langle\hat{T}_{uu}\right\rangle_{\text{ren}}$ in the Schwarzschild region just outside the collapsing shell. Our result (presented in Eq. (21)) demonstrates that option (i) above is the correct one.
@@
The geometry outside the shell (namely at $v>v_{0}$) is Schwarzschild with $M=M_{0}$, while inside the shell ($v<v_{0}$) it is Minkowski ($M=0$). That is, $M=M_{0}\Theta(v-v_{0})$, where $\Theta$ denotes the standard step function. These two different geometries are continuously matched across the shell – that is, the overall metric should be continuous (in appropriate coordinates).
@@
As already mentioned, the metric across the thin shell is in principle continuous (provided that one uses appropriate coordinates – which are themselves continuous across the shell, as we do below). In particular, $r$ is necessarily continuous at the shell, and so are the coordinates $v,\theta,\varphi$. Note, however, that if the Eddington coordinate $u$ is used globally (i.e. at both sides), the resultant metric is discontinuous. In particular, $-2g_{uv}$ is equal to $f=1-(2M_{0}/r)\Theta(v-v_{0})$, wherein the step function $\Theta(v-v_{0})$ indicates discontinuity at the shell. To achieve metric continuity, we choose our global ingoing null coordinate $\tilde{u}$ (defined outside the EH) as follows: In the Schwarzschild region outside the shell we set $\tilde{u}\equiv u$; and inside the shell we define $\tilde{u}\left(u\right)$ by simply requiring it to continuously match the aforementioned Schwarzschild’s $u$$(=\tilde{u})$. Explicitly, this continuous $\tilde{u}$ coordinate turns out to be (see (MassarParentani:1996, ; FabbriNavarroSalas:Book, ; AndersonSiahmazginClarkFabbri:2020, ))
@@
where $\hat{a}_{\omega lm}$ and $\hat{a}_{\omega lm}^{\dagger}$ are respectively the annihilation and creation operators for the $\omega lm$ mode. Our quantum state, denoted $|0\rangle$, is taken to be the “in” vacuum state, namely the vacuum state with respect to the mentioned positive-frequency modes $\Phi_{\omega lm}(x)$, obeying $\hat{a}_{\omega lm}|0\rangle=0$ (indicating no excitations originating from PNI).
@@
In view of the spherical symmetry, we choose $\theta$-splitting (LeviOriTheta:2016, ). Without loss of generality, we pick our point $x$ to be at $(u,v,\theta=0)$ and $x^{\prime}$ at $(u,v,\theta=\varepsilon)$; hence $\left\langle\hat{\Phi}^{2}\right\rangle_{\text{ren}}(x)$ effectively reduces to $\left\langle\hat{\Phi}^{2}\right\rangle_{\text{ren}}(u,v)$, $G(x,x^{\prime})$ to $G(u,v;\varepsilon)$, and $G_{DS}(x,x^{\prime})$ to $G_{DS}(r;\varepsilon)$,  [^2] In our problem the quantum state is spherically symmetric but not static (due to the collapse). However, the counterterms are independent of the quantum state and therefore respect the local staticity of the Schwarzschild metric, depending only on $r$ (and $\varepsilon$). where $r=r(u,v)$ is to be understood. The above equation (5) then reduces to
@@
Note that when computing $\left\langle\hat{\Phi}^{2}\right\rangle_{\text{ren}}^{+}(x_{\text{sh}})$ directly from Eq. (7), the most challenging piece to evaluate is $G^{+}\left(\tilde{u}_{0},v_{0};\varepsilon\right)$, as it requires solving the PDEs for the individual field modes $\Phi_{\omega lm}(x)$, subsequently summing and integrating over all mode contributions. Remarkably, this notoriously demanding quantity is entirely eliminated in Eq. (10), leaving merely the simpler, well-known counterterms.
The explicit form of the $\theta$-splitting counterterm for $\left\langle\hat{\Phi}^{2}\right\rangle$ in Schwarzschild, as a function of the splitting parameter $\varepsilon$, was computed in (LeviOriTheta:2016, ). Its general form is given in Eq. (3.10) therein, in terms of the coefficients $a,c,d$ which are specified in the equation preceding Eq. (4.2). For Schwarzschild with a general mass parameter $M$:
@@
Substituting this expression in Eq. (10) with $M=M_{0}$ for $G_{DS}^{+}$ and $M=0$ for $G_{DS}^{-}$, and setting $r=r_{0}$ in both, the shared $\varepsilon$-dependent term drops out. Our result for $\left\langle\hat{\Phi}^{2}\right\rangle_{\text{ren}}$ at the shell’s external side, in terms of $r_{0}$ (the $r$ value on the shell), is therefore:
@@
 [^4] Incidentally, when evaluating this result at $r_{0}=2M_{0}$, one obtains $\hbar/192\pi^{2}M_{0}^{2}$, which is precisely the value of $\left\langle\hat{\Phi}^{2}\right\rangle_{\text{ren}}$ at the EH of a Schwarzschild BH in the Hartle-Hawking state (Candelas:1980, ).
@@
Since the shell is parameterized by $\tilde{u}$ (along with $\theta,\varphi$), the continuity of $\Phi_{\omega lm}\left(x\right)$ across the shell implies the continuity of $\Phi_{\omega lm,\tilde{u}}$ there, allowing a natural extension of the $\left\langle\hat{\Phi}^{2}\right\rangle_{\text{ren}}$ analysis to $\left\langle\hat{T}_{\tilde{u}\tilde{u}}\right\rangle_{\text{ren}}$.
The analog of Eq. (5) for the (trace-reversed) RSET is given in Eq. (2.6) in (LeviRSET:2017, ) (which applies to any splitting direction). When reduced to the $uu$ component it reads (recalling that $g_{uu}=0$ and hence $T_{uu}$ equals its trace-reversed counterpart)
@@
We would again like to subtract the “--” version of $\left\langle\hat{T}_{uu}\right\rangle_{\text{ren}}$ from its “$+$” version, to eliminate the $G_{uu}$ terms altogether. Recall, however, that it is $\hat{\Phi}_{,\tilde{u}}$ that is continuous across the shell, not $\hat{\Phi}_{,u}$. Therefore, $G_{uu}$ is not continuous at the shell, whereas $G_{\tilde{u}\tilde{u}}$ is. This motivates transforming Eq. (14) from $u$ to $\tilde{u}$, via multiplication by $(du/d\tilde{u})^{2}$:
@@
where $G_{\tilde{u}\tilde{u}}(x,x^{\prime})\equiv\frac{1}{2}\left\langle\left\{\hat{\Phi}_{,\tilde{u}}(x),\hat{\Phi}_{,\tilde{u}}(x^{\prime})\right\}\right\rangle$ and
@@
The mentioned basic Schwarzschild’s counterterm $\tilde{L}_{uu}(r;\varepsilon)$ was presented in Sec. 1 of the supplemental material of (FluxesRN:2020, ) (setting $Q=0$). It reads
@@ Discussion.
The result (21) reveals that as the shell approaches the EH, the outflux density $\left\langle\hat{T}_{uu}\right\rangle_{\text{ren}}^{+}$ emerging from the shell vanishes quadratically with the remaining distance $r_{0}-2M_{0}$. This vanishing rate ensures the regularity of the Kruskal-based outflux density $\left\langle\hat{T}_{UU}\right\rangle_{\text{ren}}^{+}$ at the EH.
@@
Our result clarifies the origin of the Hawking outflux (at least in the collapsing null shell case). Consider the evolution of $F=4\pi r^{2}\left\langle\hat{T}_{uu}\right\rangle_{\text{ren}}$ along an outgoing null ray $u=\text{const}\equiv u_{0}\gg M_{0}$. This geodesic leaves the shell at $r=r_{0}\approx 2M_{0}$ and propagates towards FNI. Eq. (21) implies that the outflux emerging from the shell, $F^{+}\left(r_{0}\right)\equiv 4\pi r_{0}^{2}\left\langle\hat{T}_{uu}\right\rangle_{\text{ren}}^{+}\left(r_{0}\right)$, is negligibly small for such a geodesic; whereas when the geodesic approaches the weak-field region ($r\gg M_{0}$), $F$ has already reached its asymptotic value $F_{H}$. That is, the evolution of $F$ along the outgoing null ray from $F^{+}\left(u_{0}\right)\approx 0$ (at the shell) to $F_{H}$ (at FNI) occurs, gradually, as the ray traverses the strong-field region. This designates scenario (i), presented in the Introduction, as the correct physical description.
@@
Multiplying Eq. (21) by $4\pi r_{0}^{2}$, one finds that at $r_{0}^{\text{max}}=(10/3)M$ the mentioned outflux $F^{+}\left(r_{0}\right)$ reaches a maximum value of $\left(81/250000\pi\right)\hbar M^{-2}\approx 1.031\times 10^{-4}\hbar M^{-2}$. Interestingly, this value exceeds the known (Elster:1983, ) Hawking outflux in Schwarzschild, $F_{H}\approx 7.44\times 10^{-5}\hbar M^{-2}$. This suggests that perhaps the outflux $F$ evolves nonmonotonically (with $r$) along the outgoing null geodesic emanating from the shell at $r_{0}^{\text{max}}$. This should not be too surprising, as such nonmonotonic behavior of the outflux is actually observed already in the Unruh state in Schwarzschild (see Appendix).
@@
As noted in footnote [33] , our computational method inherently involves an interchange of two limits, which we have not attempted to rigorously justify. A deeper mathematical investigation of this issue of interchangeability would be desired, although such an analysis may be rather challenging. Generally speaking, the computation of renormalized quantities in BH backgrounds may typically involve several limits, including: (i) the coincidence limit $x^{\prime}\to x$, (ii) an infinite mode sum (if mode decomposition is used), and (iii) in some cases, an additional limit of approaching a desired destination point $x_{\text{dest}}$ (e.g., when a direct computation at $x_{\text{dest}}$ is problematic or challenging). Basically, the required order of limits is: first, performing the mode sum (if relevant); second, taking the coincidence limit (after counterterm subtraction); and finally, approaching the destination point (when applicable). However, in various concrete computational schemes, some of these limits are often interchanged. For instance, in the “Pragmatic Mode-sum Regularization” method (LeviOriT:2015, ; LeviOriRSET:2016, ), the mode sum and the coincidence limit are effectively interchanged. Nevertheless, results obtained through this method have been separately confirmed in multiple cases (LeviOriT:2015, ; LeviOriRSET:2016, ; LeviRSET:2017, ; LeviEilonOriMeent:2017, ; SchLanir:2018, ; KerrIH:2022, ; t-splitKerr:2024, ).
@@
It may also be interesting – and physically important – to replace the scalar field with the more realistic electromagnetic field.

### v3

@@ Abstract
We consider a minimally coupled, massless quantum scalar field $\hat{\Phi}$ propagating in the background geometry of a four-dimensional black hole formed by the collapse of a spherical thin null shell, with a Minkowski interior and a Schwarzschild exterior. The field is taken in the natural “in” vacuum state, namely, the quantum state in which no excitations arrive from past null infinity. Within the semiclassical framework, we analyze the vacuum polarization $\left\langle\hat{\Phi}^{2}\right\rangle_{\text{ren}}$ and the energy outflux density $\left\langle\hat{T}_{uu}\right\rangle_{\text{ren}}$ (where $u$ is the standard null Eddington coordinate) just outside the shell. Using the point-splitting method, we derive closed-form analytical expressions for both of these semiclassical quantities. In particular, our result for $\left\langle\hat{T}_{uu}\right\rangle_{\text{ren}}$ reveals that it vanishes like $(1-2M_{0}/r_{0})^{2}$ as the shell collapses toward the event horizon, where $M_{0}$ is the shell’s mass and $r_{0}$ is the value of the area coordinate $r$ at the evaluation point. This confirms that, along a late-time outgoing null geodesic (i.e., one that emerges from the shell very close to the event horizon and propagates toward future null infinity), the outflux gradually evolves (from virtually zero) up to its final Hawking-radiation value while the geodesic traverses the strong-field region (rather than the Hawking-radiation outflux being emitted entirely from the collapsing shell, which would lead to significant backreaction effects).
@@ Introduction.
Where does this Hawking-radiation energy come from? This has been a long-standing question (Unruh:1977, ; ChenUnruhWuYeom:2018, ). It is quite obvious that it emerges out of the strong-field region, but it is less clear exactly from what part of that region it originates. A widely accepted view (see, e.g., Ref. (Unruh:1977, ; ChenUnruhWuYeom:2018, )), to which we shall refer as option (i), is that the outgoing radiation starts from zero at the event horizon (EH) and gradually develops over a fairly broad strong-field region, until the area coordinate $r$ reaches a few times the BH mass $M$. However, there has also been an alternative viewpoint (option (ii)) suggesting that this energy outflux actually emerges directly from the collapsing body itself, and propagates in a conserved (or approximately conserved) manner through the surrounding strong-field vacuum region until it arrives FNI as Hawking radiation (see e.g. (Davies:1976, ; Gerlach:1976, ; Boulware:1976, ; KawaiMatsuoYokokura:2013, ; Ho:2015, ; KawaiYokokura:2016, ; Ho:2016, ; MaskalaniecSikorski:2024, )).
For simplicity, let us consider a spherically-symmetric collapsing object, surrounded by a vacuum region whose geometry is described by the Schwarzschild metric $ds^{2}=-f\,dt^{2}+f^{-1}dr^{2}+r^{2}d\Omega^{2}$, where $f\equiv 1-2M/r$ and $d\Omega^{2}\equiv d\theta^{2}+\sin^{2}\theta\,d\varphi^{2}$  [^1] We use general-relativistic units $c=G=1$.. We shall focus here on the domain $r>2M$ of this Schwarzschild geometry. We define the two null Eddington coordinates $u,v$ by $v\equiv t+r_{*}$ and $u\equiv t-r_{*}$, where $r_{*}\equiv r+2M\ln\left(r/2M-1\right)$ (in Minkowski, $r_{*}=r$). The energy outflux density is then represented by $\left\langle\hat{T}_{uu}\right\rangle_{\text{ren}}$, that is, the outgoing null component of the renormalized stress-energy tensor (RSET) $\left\langle\hat{T}_{\mu\nu}\right\rangle_{\text{ren}}(x)$.
@@
Evidently, this quantity diverges on approaching the EH, which would potentially lead to drastic backreaction effects – such as entirely halting the collapse process (Gerlach:1976, ; Boulware:1976, ; KawaiMatsuoYokokura:2013, ; Mersini:2014(HH), ; Mersini:2014(Unruh), ; Ho:2015, ; KawaiYokokura:2016, ; Ho:2016, ; MaskalaniecSikorski:2024, ) and/or preventing the EH formation (for various scenarios of “horizon avoidance” in semiclassical collapse see, e.g., (Gerlach:1976, ; KawaiMatsuoYokokura:2013, ; Ho:2015, ; KawaiYokokura:2016, ; Ho:2016, ; MaskalaniecSikorski:2024, ; Mersini:2014(HH), ; Mersini:2014(Unruh), ; BMT:2016, ; BMT:2017a, ; BMT:2017b, )).
@@
We hereafter model the collapse process by a thin null massive shell, with an empty, flat interior (more details below). For simplicity we consider a minimally-coupled massless scalar field $\Phi(x)$. Several authors (see, e.g., Refs. (AndersonSiahmazginClarkFabbri:2020, ; SiahmazgiAndersonClarkFabbri:2021, ; SiahmazgiAndersonFabbri:2025, ) and references within) have already elaborated on this model, but so far, to our knowledge, a concrete result for $\left\langle\hat{T}_{uu}\right\rangle_{\text{ren}}$ in the strong-field region has not been obtained. We shall derive a simple, explicit, analytical expression for $\left\langle\hat{T}_{uu}\right\rangle_{\text{ren}}$ in the Schwarzschild region just outside the collapsing shell. Our result (presented in Eq. (21)) demonstrates that option (i) above is the correct one.
@@
The geometry outside the shell (namely at $v>v_{0}$) is Schwarzschild with $M=M_{0}$, while inside the shell ($v<v_{0}$) it is Minkowski ($M=0$), that is, $M=M_{0}\Theta(v-v_{0})$, where $\Theta$ denotes the standard step function. These two different geometries are continuously matched across the shell – that is, the overall metric should be continuous (in appropriate coordinates).
@@
As already mentioned, the metric across the thin shell is in principle continuous (provided that one uses appropriate coordinates – which are themselves continuous across the shell, as we do below). In particular, $r$ is necessarily continuous at the shell, and so are the coordinates $v,\theta,\varphi$. Note, however, that if the Eddington coordinate $u$ is used globally (i.e. at both sides), the resultant metric is discontinuous. In particular, $-2g_{uv}$ is equal to $f=1-(2M_{0}/r)\Theta(v-v_{0})$, wherein the step function $\Theta(v-v_{0})$ indicates discontinuity at the shell. To achieve metric continuity, we introduce a new global null coordinate $\tilde{u}$ (defined outside the EH) as follows: in the Schwarzschild region outside the shell we set $\tilde{u}\equiv u$; and inside the shell we define $\tilde{u}\left(u\right)$ by simply requiring it to continuously match the aforementioned Schwarzschild’s $u$$(=\tilde{u})$. Explicitly, this continuous $\tilde{u}$ coordinate turns out to be (see (MassarParentani:1996, ; FabbriNavarroSalas:Book, ; AndersonSiahmazginClarkFabbri:2020, ))
@@
where $\hat{a}_{\omega lm}$ and $\hat{a}_{\omega lm}^{\dagger}$ are the annihilation and creation operators for the $\omega lm$ mode, respectively. Our quantum state, denoted $|0\rangle$, is taken to be the “in” vacuum state, namely the vacuum state with respect to the mentioned positive-frequency modes $\Phi_{\omega lm}(x)$, obeying $\hat{a}_{\omega lm}|0\rangle=0$ (indicating no excitations originating from PNI).
@@
In view of the spherical symmetry, we choose $\theta$-splitting (LeviOriTheta:2016, ). Without loss of generality, we pick our point $x$ to be at $(u,v,\theta=0)$ and $x^{\prime}$ at $(u,v,\theta=\varepsilon)$; hence $\left\langle\hat{\Phi}^{2}\right\rangle_{\text{ren}}(x)$ effectively reduces to $\left\langle\hat{\Phi}^{2}\right\rangle_{\text{ren}}(u,v)$, $G(x,x^{\prime})$ to $G(u,v;\varepsilon)$, and $G_{DS}(x,x^{\prime})$ to $G_{DS}(r;\varepsilon)$,  [^2] In our problem the quantum state is spherically symmetric but not static (due to the collapse). However, the counterterms are independent of the quantum state and therefore respect the local staticity of the Schwarzschild metric, depending only on $r$ (and $\varepsilon$). where $r=r(u,v)$ is to be understood. Equation (5) then reduces to
@@
Note that when computing $\left\langle\hat{\Phi}^{2}\right\rangle_{\text{ren}}^{+}(x_{\text{sh}})$ directly from Eq. (7), the most challenging piece to evaluate is $G^{+}\left(\tilde{u}_{0},v_{0};\varepsilon\right)$, as it requires solving the partial differential equations for the individual field modes $\Phi_{\omega lm}(x)$, subsequently summing and integrating over all mode contributions. Remarkably, this notoriously demanding quantity is entirely eliminated in Eq. (10), leaving merely the simpler, well-known counterterms.
The explicit form of the $\theta$-splitting counterterm for $\left\langle\hat{\Phi}^{2}\right\rangle$ in Schwarzschild, as a function of the splitting parameter $\varepsilon$, was computed in (LeviOriTheta:2016, ). Its general form is given in Eq. (3.10) therein, in terms of the coefficients $a,c,d$ which are specified in the equation preceding Eq. (4.2) therein. For Schwarzschild with a general mass parameter $M$,
@@
Substituting this expression in Eq. (10) with $M=M_{0}$ for $G_{DS}^{+}$ and $M=0$ for $G_{DS}^{-}$, and setting $r=r_{0}$ in both, the shared $\varepsilon$-dependent term drops out. Our result for $\left\langle\hat{\Phi}^{2}\right\rangle_{\text{ren}}$ at the shell’s external side, in terms of $r_{0}$ (the $r$ value on the shell), is therefore  [^4] Incidentally, when evaluating this result at $r_{0}=2M_{0}$, one obtains $\hbar/192\pi^{2}M_{0}^{2}$, which is precisely the value of $\left\langle\hat{\Phi}^{2}\right\rangle_{\text{ren}}$ at the EH of a Schwarzschild BH in the Hartle-Hawking state (Candelas:1980, ).
@@
Since the shell is parameterized by $\tilde{u}$ (along with $\theta,\varphi$), the continuity of $\Phi_{\omega lm}\left(x\right)$ across the shell implies the continuity of $\Phi_{\omega lm,\tilde{u}}$ there, allowing for a natural extension of the $\left\langle\hat{\Phi}^{2}\right\rangle_{\text{ren}}$ analysis to $\left\langle\hat{T}_{\tilde{u}\tilde{u}}\right\rangle_{\text{ren}}$ (and hence also to $\left\langle\hat{T}_{uu}\right\rangle_{\text{ren}}$).
The analog of Eq. (5) for the (trace-reversed  [^5] The trace reverse of $T_{\alpha\beta}$ is defined as $T_{\alpha\beta}-g_{\alpha\beta}T_{\,\,\mu}^{\mu}/2$.) RSET is given in Eq. (2.6) in (LeviRSET:2017, ) (which applies to any splitting direction). When reduced to the $uu$ component it reads (recalling that $g_{uu}=0$ and hence $T_{uu}$ equals its trace-reversed counterpart)
@@
We would again like to subtract the “--” version of $\left\langle\hat{T}_{uu}\right\rangle_{\text{ren}}$ from its “$+$” version, to eliminate the $G_{uu}$ terms altogether. Recall, however, that it is $\hat{\Phi}_{,\tilde{u}}$ that is continuous across the shell, not $\hat{\Phi}_{,u}$. Therefore, $G_{uu}$ is not continuous at the shell, whereas $G_{\tilde{u}\tilde{u}}$ is, where $G_{\tilde{u}\tilde{u}}(x,x^{\prime})\equiv\frac{1}{2}\left\langle\left\{\hat{\Phi}_{,\tilde{u}}(x),\hat{\Phi}_{,\tilde{u}}(x^{\prime})\right\}\right\rangle$. This motivates transforming Eq. (14) from $u$ to $\tilde{u}$, via multiplication by $(du/d\tilde{u})^{2}$:
@@
where
@@
The mentioned basic Schwarzschild’s counterterm $\tilde{L}_{uu}(r;\varepsilon)$ was presented in Sec. I of the Supplemental Material of (FluxesRN:2020, ) (setting $Q=0$). It reads
@@ Discussion.
The result (21) reveals that as the shell approaches the EH, the outflux density $\left\langle\hat{T}_{uu}\right\rangle_{\text{ren}}^{+}$ emerging from the shell vanishes quadratically with the remaining distance $r_{0}-2M_{0}$. This vanishing rate ensures the regularity of the Kruskal-based outflux density $\left\langle\hat{T}_{UU}\right\rangle_{\text{ren}}^{+}$ at the EH. Since the RSET magnitude is typically of order $\hbar/M^{2}\lll 1$, this regularity actually implies that semiclassical backreaction cannot prevent (or significantly affect) the horizon formation.
@@
Our result clarifies the origin of the Hawking outflux (at least in the collapsing null shell case). Consider the evolution of $F=4\pi r^{2}\left\langle\hat{T}_{uu}\right\rangle_{\text{ren}}$ along an outgoing null ray $u=\text{const}\equiv u_{0}\gg M_{0}$. This geodesic leaves the shell at $r=r_{0}\approx 2M_{0}$ and propagates toward FNI. Equation (21) implies that the outflux emerging from the shell, $F^{+}\left(r_{0}\right)\equiv 4\pi r_{0}^{2}\left\langle\hat{T}_{uu}\right\rangle_{\text{ren}}^{+}\left(r_{0}\right)$, is negligibly small for such a geodesic; whereas when the geodesic approaches the weak-field region ($r\gg M_{0}$), $F$ has already reached its asymptotic value $F_{H}$. That is, the evolution of $F$ along the outgoing null ray from $F^{+}\left(u_{0}\right)\approx 0$ (at the shell) to $F_{H}$ (at FNI) occurs, gradually, as the ray traverses the strong-field region. This designates scenario (i), presented in the Introduction, as the correct physical description.
@@
Multiplying Eq. (21) by $4\pi r_{0}^{2}$, one finds that at $r_{0}^{\text{max}}=(10/3)M$ the mentioned outflux $F^{+}\left(r_{0}\right)$ reaches a maximum value of $\left(81/250000\pi\right)\hbar M^{-2}\approx 1.031\times 10^{-4}\hbar M^{-2}$. Interestingly, this value exceeds the known (Elster:1983, ) Hawking outflux in Schwarzschild, $F_{H}\approx 7.44\times 10^{-5}\hbar M^{-2}$. This suggests that perhaps the outflux $F$ evolves nonmonotonically (with $r$) along the outgoing null geodesic emanating from the shell at $r_{0}^{\text{max}}$. This should not be too surprising, as such nonmonotonic behavior of the outflux is actually observed already in the Unruh state in Schwarzschild (see (Sup, )).
@@
As noted in footnote [413], our computational method inherently involves an interchange of two limits, which we have not attempted to rigorously justify. A deeper mathematical investigation of this issue of interchangeability would be desired, although such an analysis may be rather challenging. Generally speaking, the computation of renormalized quantities in BH backgrounds may typically involve several limits, including: (i) the coincidence limit $x^{\prime}\to x$, (ii) an infinite mode sum (if mode decomposition is used), and (iii) in some cases, an additional limit of approaching a desired destination point $x_{\text{dest}}$ (e.g., when a direct computation at $x_{\text{dest}}$ is problematic or challenging). Basically, the required order of limits is as follows: first, performing the mode sum (if relevant); second, taking the coincidence limit (after counterterm subtraction); and finally, approaching the destination point (when applicable). However, in various concrete computational schemes, some of these limits are often interchanged. For instance, in the “Pragmatic Mode-sum Regularization” method (LeviOriT:2015, ; LeviOriRSET:2016, ), the mode sum and the coincidence limit are effectively interchanged. Nevertheless, results obtained through this method have been separately confirmed in multiple cases (LeviOriT:2015, ; LeviOriRSET:2016, ; LeviRSET:2017, ; LeviEilonOriMeent:2017, ; SchLanir:2018, ; KerrIH:2022, ; t-splitKerr:2024, ).
@@
In the analogous 2D case the RSET is known explicitly (Hiscock:1981, )(FabbriNavarroSalas:Book, ), and one can easily see that $\left\langle T_{\tilde{u}\tilde{u}}\right\rangle$ actually vanishes at both sides of the collapsing null shell – in contrast with the discontinuity of $\left\langle T_{\tilde{u}\tilde{u}}\right\rangle$ at the shell in 4D. This sharp distinction may be rooted in two obvious differences between the two cases. First and most importantly, in the 2D case $\left\langle T_{\alpha}^{\alpha}\right\rangle$ is known (from the trace anomaly (Davies:1977, )) to be proportional to the background’s Ricci scalar – which is bounded at the shell. From this boundedness one can easily show (using energy-momentum conservation) that $\left\langle T_{\tilde{u}\tilde{u}}\right\rangle$ must be continuous at the null shell. Such a constraint does not apply in 4D, primarily because $\left\langle T_{\alpha\beta}\right\rangle$ now involves one additional degree of freedom (related to the angular RSET components). Second, recall that in 4D our minimally-coupled field $\Phi(x)$ is non-conformal, hence $\left\langle T_{\alpha}^{\alpha}\right\rangle$ now includes an extra term $\propto\boxempty\langle\Phi^{2}\rangle$ (Group:2018, ) – which in fact involves $\delta(v-v_{0})$. In view of this delta function in $\left\langle T_{\alpha}^{\alpha}\right\rangle$ in 4D, a simple analysis (involving energy-momentum conservation) leads to the anticipation for a step function in $\left\langle T_{\tilde{u}\tilde{u}}\right\rangle$ at the shell – which was indeed found here explicitly.
It may be interesting – and physically important – to extend the analysis to the more realistic case of an electromagnetic field.